OpenStudy (anonymous):

Given f(x) = the quantity of x plus 7, divided by 5, solve for f-1(3).

OpenStudy (the_fizicx99):

Inverse function xD Start by replacing \(\ \sf f(x)\) with \(\ \sf y\) \(\ \sf f(x) = \dfrac{x + 7}{5} \rightarrow y = \dfrac{x + 7}{5} \) Now \(\ \sf exchange \) y with your x so \(\ \sf y = \dfrac{x + 7}{5} \) becomes \(\ \sf x = \dfrac{y + 7}{5} \) and now you want to solve for y. Multiply both sides by 5 and subtract 7 from both sides. \(\ \sf x = \dfrac{y + 7}{5} \longrightarrow x * 5 ~ or ~ 5x\) \(\ \sf x = \dfrac{y + 7}{5} \longrightarrow 5x - 7 = y\) Now Simply repleace y with f(x)\(^{-1}\) \(\ \sf \Large 5x - 7 = y \longrightarrow 5x - 7 = f(x)^{-1} \) Now plug in 3 for x.

OpenStudy (the_fizicx99):

\(\ \sf \Large f(\color{orange}{x})^{-1} = 5\color{orange}{x} - 7 \) \(\ \sf \color{orange}{x = 3} \) \(\ \sf \Large f(\color{orange}{3})^{-1} = 5(\color{orange}{3}) - 7 \rightarrow 15 - 7 = 8\) f(3)^-1 = 8

OpenStudy (anonymous):

@Kimberly_PR